OpenStudy (anonymous):

A scientist is studying the growth of a particular species of plant. He writes the following equation to show the height of the plant f(n), in cm, after n days: f(n) = 12(1.03)n Part A: When the scientist concluded his study, the height of the plant was approximately 16.13 cm. What is a reasonable domain to plot the growth function? (4 points) Part B: What does the y-intercept of the graph of the function f(n) represent? (2 points) Part C: What is the average rate of change of the function f(n) from n = 3 to n = 10, and what does it represent? (4 points

OpenStudy (anonymous):

@dan815

OpenStudy (dan815):

hey im here

OpenStudy (dan815):

you are missing exponents on this equation f(n) = 12(1.03)n

OpenStudy (anonymous):

What do you mean?

OpenStudy (dan815):

i think your equation is wrong

OpenStudy (anonymous):

it is correctly wrriten

OpenStudy (dan815):

are you absolutely sure, because, they do not ask growth rate for non exponential equations

OpenStudy (anonymous):

f(n) = 12(1.03)^n

OpenStudy (dan815):

attaboy

OpenStudy (anonymous):

(:

OpenStudy (dan815):

mhm

OpenStudy (anonymous):

Ok so what is next?

OpenStudy (dan815):

now lets do part A what height did the plant end at

OpenStudy (anonymous):

16.13cm

OpenStudy (dan815):

16.13 = 12(1.03)^n so we have find out what day this was

OpenStudy (dan815):

this will tell us approximately how high you want to make our domain on the n axis

OpenStudy (dan815):